OpenStudy (anonymous):

Reduce this Algebraic fraction: 4x^3/3 * (9/x)^3 = ____ The answer I got was 972 but I am not so sure >_<

OpenStudy (anonymous):

(4x^3/3) * (9/x)^3 = ( 4x^3 / 3 )( 729 / x^3 ) = 2916x^3 / 3x^3 = 972

OpenStudy (anonymous):

Thank you so much ^-^